# Normal forms for ordinary differential operators, I

### v3

@@ 2.1 Preliminary statements from the Schur theory for the ring $\hat{D}_{1}^{sym}$
The proof will follow the proof of Prop. 7.2 in [46]. $P$ is regular, since $\sigma(P)=\partial^{k}$. Recall that there exist $S$ such that $S_{[0]}=1$, $S_{[i]}=0$ for ($0<i<k$) and each slice can be found from the system
@@
Therefore, the homogeneous decomposition of $S$ has no homogeneous terms of order $-1$, and $S_{1}=0$.
Lemma 2.4.
Let $P\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ be a homogeneous operator commuting with $\partial^{k}$ and $-k<\mathop{\rm\bf ord}(P)=l<0$ (if $\mathop{\rm\bf ord}(P)\leq-k$ then $P=0$ by proposition 2.1, item 2). Then
$P=(\sum_{j=0}^{k-1}c_{j}A_{j})\int^{-l},$
where $c_{j}\in\tilde{K}$ satisfy the following conditions:
or, equivalently,
$\sum_{j=0}^{k-1}c_{j}\xi^{j(q-1)}=0\quad\mbox{for\quad}q=1,\ldots,-l.$
Vice versa, any such operator $P$ commutes with $\partial^{k}$.
Proof.
All assertions of lemma can be deduced from the proof of [46, Prop.7.1]. Another proof is as follows.
hence $\sum_{j=0}^{k-1}c_{j}\xi^{j(q-1)}=0$ for $q=1,\ldots,-l$.
The same calculation proves the last assertion. ∎
@@ 2.2 Basic formulae in $\hat{D}_{1}^{sym}$
First we define a series of operators $B_{i}$: $B_{1}:=\delta,B_{2}:=x\delta\partial,\ldots,B_{n}:=\frac{1}{(n-1)!}x^{n-1}\delta\partial^{n-1}$, and $B_{j}:=0$ for any $j\leq 0$. Define $\Gamma_{i}=(x\partial)^{i}$ for $i\geq 0$, and for $i<0,\Gamma_{i}=0$. For convenience, we introduce also a new notation: for any integer $n$ we set $D^{n}=\partial^{n}$ if $n\geq 0$ and $D^{n}=\int^{-n}$ otherwise. Obviously, we have $\partial\delta=\delta x=0$ and $\mathop{\rm\bf ord}(B_{j})=0$ for any $j\in\mathbb{N}$.
@@
    $D^{u}B_{j}=B_{j-u}D^{u}$. In particular, $D^{u}B_{j}=0$ if $u>0$ and $j-u\leq 0$ or $u<0$ and $j-u\leq-u$,
where we assume in all these formulae that $10^{0}:=$ 1.
@@
$\int^{m}\cdot\partial^{m}=\int\cdot\int^{m-1}\cdot\partial^{m-1}\cdot\partial=\int\cdot(1-\sum_{k=1}^{m-1}B_{k})\cdot\partial$
$=\int\cdot(1-\sum_{k=1}^{m-2}B_{k})\cdot\partial+\int\cdot(B_{m-1})\cdot\partial=\int\cdot\int^{m-2}\cdot\partial^{m-2}\cdot\partial+\int\cdot(B_{m-1})\cdot\partial$
$=1-\sum_{k=1}^{m-1}B_{k}-\int(\frac{1}{(m-2)!}x^{m-2}\delta\partial^{m-2})\partial=1-\sum_{k=1}^{m-1}B_{k}-B_{m}.$
4. Note that the equality $\int x=x\int-\int^{2}$ hold iff $\int x\partial^{q}=(x\int-\int^{2})\partial^{q}$ hold for any $q\geq 0$ (as it follows from definition of the ring $\hat{D}_{1}^{sym}$). Take $q=2$. Then we have
@@
So, $\int x\partial^{2}=(x\int-\int^{2})\partial^{2}$ and therefore $\int x=x\int-\int^{2}$. The second formula follows immediately by induction (note that each summand is homogeneous of order $(-u+l)$ and therefore the total sum is well defined for any series $f(x)$).
@@
$A_{i}(x\partial)^{j}=(x\partial)A_{i}(x\partial)^{j-1}=\cdots=(x\partial)^{j}A_{i}$
@@ 2.3 Homogeneous canonical polynomials
Obviously, of all coefficients are equal to zero, then $H=0$. Now assume the converse:
@@
where $f_{d_{i},i;r}^{\prime}\in\tilde{K}$ are some new coefficients, but $f_{d_{i},i;r}^{\prime}=f_{d_{i},i;r}$ for all $i$.
@@
for some new $^{\prime}\in{\mbox{\dbl N}}$ ℕ, $g_{j;r}^{\prime}\in\tilde{K}$, but with the same coefficients $f_{l,i;r}^{\prime}$. Indeed, if $r\geq 0$, then $D^{r}D^{-r}=1$ and therefore all coefficients of $H^{\prime}$ are the same. If $r<0$, then by lemma 2.5 $D^{r}D^{-r}=1-\sum_{i=1}^{r}B_{i}$, and by the same lemma any product
$[\sum_{0\leq i<k}\sum_{0\leq l\leq d_{i}}f_{l,i;r}^{\prime}\xi^{-il}A_{k;i}\Gamma_{l}+\sum_{0<j\leq N}g_{j;r}B_{j}]B_{i}$
@@
It’s easy to see that two forms of $H$ are one to one correspondence to each other, and therefore the G-form is also uniquely defined. Besides, the definitions of $Sdeg_{A}$, $Sdeg_{B}$ does not depend on the form, i.e. $Sdeg_{A}(H)$ is again the maximal $d_{i}$.
@@
([34, Ch.2, item 10 Th.8]) Suppose $L,f\in\tilde{K}[t]$ are non-zero polynomials with $\deg f=r$, $\lambda\in\tilde{K}$, and $q$ is the multiplicity of $t-\lambda$ in the polynomial $L(t)$ (if $L(\lambda)\neq 0$, then $q=0$). Then the ODE
@@
    If $Sdeg_{A}(M)=-\infty$ then $Sdeg_{A}(H)$ is either 00 or $\infty$- ∞ and $Sdeg_{B}(H)=Sdeg_{B}(M)$.
@@
$H^{(k)}+\dots+kH^{\prime}\tilde{\partial}^{k-1}=M$ (2.5)
The eigen-polynomial of this ODE is $L(t):=(t-\tilde{\partial})^{k}-\tilde{\partial}^{k}$. Any solution $H$ of this equation looks like $H=H_{0}+\sum_{i=0}^{k-1}c_{i}\tilde{A}_{k,i}$, where $c_{i}\in\tilde{K}((\tilde{\partial}))$ and $H_{0}$ is a special solution. Since $M$ is a HCP, it can be represented as a linear combination of quasi-polynomials in the ring $\tilde{R}$, which are homogeneous with respect to the order $\mathop{\rm\bf ord}$. Therefore, there is a partial solution equal to a sum of partial solutions of similar equations with monomial right hand side $M$.
2. We have two possible case of such monomials:
@@
    $M=f_{i}(x)\tilde{A}_{i}$. Notice that $\tilde{A}_{i}=e^{(\xi^{i}-1)x\tilde{\partial}}$ and $(\xi^{i}-1)\tilde{\partial}$ is the root of $L(t)$ with multiplicity 1. Then according to Lemma 2.8, such ODE has a special solution $H_{0}=xg_{i}(x)\tilde{A}_{i}$ (with $deg_{x}(f_{i})=deg_{x}(g_{i})$).
@@
    $M=b_{j}\tilde{B}_{j}$, where $b_{j}\in\tilde{K}$. Notice that $\tilde{B}_{j}=cx^{j-1}\tilde{\partial}^{j-1}\tilde{\delta}=cx^{j-1}\tilde{\partial}^{j-1}e^{-x\tilde{\partial}}$, and $-\tilde{\partial}$ is not a root of $L(t)$. Then according to Lemma 2.8, we know the special solution of this ODE will be in the form
    $H_{0}=h_{j}(x)\tilde{\delta}$
@@
3. Now we have: $H=H_{0}+\sum_{i=0}^{k-1}c_{i}\tilde{A}_{k,i}$, where $c_{i}\in\tilde{K}((\tilde{\partial}))$ and $H_{0}$ is a HCP. Moreover, if $Sdeg_{A}(M)\neq-\infty$, then $Sdeg_{A}(H_{0})=Sdeg_{A}(M)+1$ and $Sdeg_{B}(H_{0})=Sdeg_{B}(M)$. And if $Sdeg_{A}(M)=-\infty$, then $Sdeg_{A}(H_{0})=-\infty$ and $Sdeg_{B}(H_{0})=Sdeg_{B}(M)$. Since our original solution $H\in\hat{D}_{1}^{sym}$, we have $\mathop{\rm\bf ord}(H)<\infty$ and therefore $\mathop{\rm\bf ord}(c_{i})<\infty$ for all $i$, i.e. $c_{i}\in\tilde{K}[\tilde{\partial}^{-1},\tilde{\partial}]$ and $H$ is a HCPC.
Writing this solution in a canonical form and replacing $\tilde{\partial}$ by $D$, we get, as in the proof of propostion 2.1, that $H$ is a HCPC. Since $Sdeg_{A}(\sum_{i=0}^{k-1}c_{i}{A}_{k,i})=0$, we get our assertion. Thus we complete the proof.
@@
The condition about existence of $H$ from lemma is essential: for example $[\partial^{4},H]=B_{10}\partial^{4}$ doesn’t have any HCPC solution. But we will see if $M$ doesn’t contain $B_{j}$ such $H$ must exists, see lemma 2.11.
@@
    $Sdeg_{A}(T)\leq Sdeg_{A}(H)+Sdeg_{A}(M)$ (here we assume $-\infty+n=-\infty$).
@@
Consider the following 4 cases (below we assume in all formulae that $10^{0}:=$ 1):
@@
(here $\delta_{i}^{j-u}$ is the Kronecker delta), because if $i=j-u$, then $i-1+u+v=j-1+v\geq 0$ and $B_{i}D^{u+v}\neq 0$.
2. $H=b_{i}B_{i}D^{u}$, where $i\geq 1-u$ if $u<0$, $M=a_{l,m}\Gamma_{m}A_{j}D^{v}$, $b_{i},a_{l,m}\in\tilde{K}$. By Lemma 2.5 item 12 we know $B_{i}D^{u}=D^{u}B_{i+u}$, hence $i\geq 1-u$, otherwise $H=0$. Then we have
@@
And by item 1 we know $B_{i}B_{j}=\delta^{i}_{j}B_{i}$, hence we know if $-u-v+1\leq i\leq-u$, then
@@
But we already assume $i\geq 1-u$, hence $B_{i}D^{u}D^{v}=B_{i}D^{u+v}$, again by Lemma 2.5 item 9, we know it is 0 when $i-1+u+v<0$.
@@
where $\lambda=a_{l,m}b_{i}\xi^{j(i-v-1)}(i-v-1)^{m}$. This is because when $i-v<1$ we have $D^{v}B_{i}=B_{i-v}D^{v}=0$ (c.f. Lemma 2.5 item 9 and by assumptions for $r\leq 0$, $B_{r}=0$), so that
@@
The last equality holds obviously when $v\geq 0$ or $u\leq 0$. When $v<0,u>0$, notice that $i>0$, so that $i-v>-v$, hence for any $-u-v+1\leq s\leq-v$, we have $B_{i-v}B_{s}=\delta_{s}^{i-v}=0$ by Lemma 2.5 item 5. Hence
@@ 1. Assume
Since $[\partial^{k},A_{i}]=0$, we know by lemma 2.10 that
$[\partial^{k},H_{i;m-k}A_{i}D^{m-k}]=M_{i;m}A_{i}D^{m}\qquad[\mod B_{j}],$ (2.10)
where $[\mod B_{j}]$ denote denote terms containing $B_{j}$s. Now assume
@@
Since $[\partial^{k},h_{0,i;m-k}A_{i}D^{m-k}]$ contains only terms with $B_{j}$, we can ignore terms with $h_{0,i;m-k}$ by looking for a solution. Assume
@@
Using Lemma 2.10, we can directly calculate $[\partial^{k},H_{i;m-k}A_{i}D^{m-k}]$. Then we get the linear system for the unknown coefficients ${h_{1,i;m-k},\dots,h_{t_{i},i;m-k}}$:
$\begin{pmatrix}k&*&\dots&*\\ 0&2k&\dots&*\\ \dots&\dots&\dots&\dots\\ 0&0&\dots&t_{i}k\end{pmatrix}\begin{pmatrix}h_{1,i;m-k}\\ h_{2,i;m-k}\\ \dots\\ h_{t_{i},i;m-k}\end{pmatrix}=\begin{pmatrix}m_{0,i;m}\\ m_{2,i;m}\\ \dots\\ m_{t_{i}-1,i;m}\end{pmatrix}$ (2.11)
@@
$[\partial^{k},\tilde{H}]=M\quad[\mod B_{j}]$
Now assume with
@@
with $H=\tilde{H}+\bar{H}$. The equation $[\partial^{k},H]=M$ becomes
$[\partial^{k},\bar{H}]=M-[\partial^{k},\tilde{H}]$ (2.12)
@@
    $m\geq k$, in such case no $B_{j}$ can appear on the right hand side of (2.10). So we can simply put $h_{0,0;m-k},\dots,h_{0,k-1;m-k}$ all 0. We will get $[\partial^{k},H]=M$.
@@
    where $b_{1-m},\dots,b_{k-m}$ are already known on the right hand side. Collecting this linear system, we find we have $k$ variables and $k$ equations, the coefficient matrix is a Vandermonde matrix (hence always of full rank). Thus we can always solve the equation and find the coefficients $h_{0,0;m-k},\dots,h_{0,k-1;m-k}$. After that we get $[\partial^{k},\tilde{H}+\bar{H}]=M$.
@@
    $0<m<k$, in this case the same arguments as in case b) work, and we omit the details. We only need to notice that here we have $k$ variables but $(k-m)$ equations, and the coefficient matrix becomes the full rank sub-matrix of the Vandermonde matrix. So the solution exists but might be not unique.

### v4

@@ 2.1 Preliminary statements from the Schur theory for the ring $\hat{D}_{1}^{sym}$
The proof will follow the proof of Prop. 7.2 in [49]. $P$ is regular, since $\sigma(P)=\partial^{k}$. Recall that there exists $S$ such that $S_{[0]}=1$, $S_{[i]}=0$ for $0<i<k$ and each slice can be found from the system
@@
Therefore, the homogeneous decomposition of $S$ has no homogeneous terms of order 1-1, and $S_{-1}=0$.
@@ 2.2 Basic formulae in $\hat{D}_{1}^{sym}$
First we define a series of operators $B_{i}$: $B_{1}:=\delta,B_{2}:=x\delta\partial,\ldots,B_{n}:=\frac{1}{(n-1)!}x^{n-1}\delta\partial^{n-1}$, and $B_{j}:=0$ for any $j\leq 0$. Define $\Gamma_{i}=(x\partial)^{i}$ for $i\geq 0$, and for $i<0,\Gamma_{i}=0$. For convenience, we introduce also a new notation: for any integer $n$ we set $D^{n}=\partial^{n}$ if $n\geq 0$ and $D^{n}=\int^{-n}$ otherwise. Obviously, we have $\partial\delta=\delta x=0$ and $\mathop{\rm\bf ord}(B_{j})=0$ for any $j\in\mathbb{N}$. From this observation and definition of $\delta$ we also get $\delta\delta=\delta$.
@@
    $D^{u}B_{j}=B_{j-u}D^{u}$. In particular, $D^{u}B_{j}=B_{j-u}D^{u}=0$ if $u>0$ and $j-u\leq 0$ or $u<0$ and $j\leq 0$.
where we assume in all these formulae that $0^{0}:=1$.
@@
$\int^{m}\partial^{m}=\int\int^{m-1}\partial^{m-1}\partial=\int(1-\sum_{k=1}^{m-1}B_{k})\partial\\ =\int(1-\sum_{k=1}^{m-2}B_{k})\partial-\int(B_{m-1})\partial=\int\int^{m-2}\partial^{m-2}\partial-\int(B_{m-1})\partial\\ =1-\sum_{k=1}^{m-1}B_{k}-\int(\frac{1}{(m-2)!}x^{m-2}\delta\partial^{m-2})\partial=1-\sum_{k=1}^{m-1}B_{k}-B_{m}.$
4. Note that the equality $\int x=x\int-\int^{2}$ holds iff $\int x\partial^{q}=(x\int-\int^{2})\partial^{q}$ holds for any $q\geq 0$ (as it follows from definition of the ring $\hat{D}_{1}^{sym}$). Take $q=2$. Then we have
@@
So, $\int x\partial^{2}=(x\int-\int^{2})\partial^{2}$ and therefore $\int x=x\int-\int^{2}$. The second formula follows immediately by induction for $f(x)=x^{l}$, and for generic series $f(x)=\sum_{l}c_{l}x^{l}$ the left and right hand side of the formula are just sums of homogeneous operators $\int^{u}c_{l}x^{l}$ (note that each such summand is homogeneous of order (ul)(-u-l) and therefore the total sum is well defined for any series f(x)f(x)) [^6] This formula coincides with analogous formula for the classical ring of pseudo-differential operators, cf. e.g. [50, Def. 4.1].
@@
$A_{i}(x\partial)^{j}=(x\partial)A_{i}(x\partial)^{j-1}=s=(x\partial)^{j}A_{i}$
@@ 8. First note that
The next Lemma can be deduced from the proof of [49, Prop.7.1]. Another proof, by using our calculations above is as follows:
Lemma 2.5.
Let $P\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ be a homogeneous operator commuting with $\partial^{k}$ and $k<\mathop{\rm\bf ord}(P)=l<0$ (if $\mathop{\rm\bf ord}(P)\leq-k$ then $P=0$ by proposition 2.1, item 2). Then
$P=(\sum_{j=0}^{k-1}c_{j}A_{j})\int^{-l},$
where $c_{j}\in\tilde{K}$ satisfy the following conditions:
or, equivalently,
$\sum_{j=0}^{k-1}c_{j}\xi^{j(q-1)}=0\quad\mbox{for\qquad}q=1,\ldots,-l.$
Vice versa, any such operator $P$ commutes with $\partial^{k}$.
Proof.
hence $\sum_{j=0}^{k-1}c_{j}\xi^{j(q-1)}=0$ for $q=1,\ldots,-l$.
The same calculation proves the last assertion. ∎
@@ 2.3 Homogeneous canonical polynomials
Obviously, if all coefficients are equal to zero, then $H=0$. Now assume the converse:
@@
where $f_{l,i;r}^{\prime}\in\tilde{K}$ are some new coefficients, but $f_{d_{i},i;r}^{\prime}=f_{d_{i},i;r}$ for all $i$.
@@
[…]
$\sum_{0\leq i<k}\sum_{0\leq l\leq d_{i}}f_{l,i;r}^{\prime}\xi^{-il}A_{k;i}\Gamma_{l}+\sum_{0<j\leq N}g_{j;r}B_{j}]B_{i}$
@@
It’s easy to see that two forms of $H$ are one to one correspondence to each other, and therefore the G-form is also uniquely defined. Besides, the definitions of $Sdeg_{A}$, $Sdeg_{B}$ don’t depend on the form, i.e. $Sdeg_{A}(H)$ is again the maximal $d_{i}$.
@@
([34, Ch.2, item 10 Th.8]) Let $F$ be a field of characteristic 0. Suppose $L,f\in F[t]$ are non-zero polynomials with $\deg f=r$, $\lambda\in F$, and $q$ is the multiplicity of $t-\lambda$ in the polynomial $L(t)$ (if $L(\lambda)\neq 0$, then $q=0$). Then the ODE
@@
    If $Sdeg_{A}(M)=$-$\infty$ then $Sdeg_{A}(H)$ is either 0 or -$\infty$ and $Sdeg_{B}(H)=Sdeg_{B}(M)$.
@@
1. Consider the HCP decomposition of $M$: $M=\sum_{i=l_{1}}^{l_{2}}M_{i}$, where $\mathop{\rm\bf ord}(M_{i})=i$, and the HCP decomposition of $H$. One can easily see that
$\partial^{k},H_{i-k}]=M_{i}$
holds for any $i$. Since $Sdeg_{A},Sdeg_{B}$ of $H,M$ are defined as the maximum values of $Sdeg_{A},Sdeg_{B}$ on $H_{i},M_{j}$, we can assume that $H,M$ are HCPs.
@@
$H^{(k)}+\dots+kH^{\prime}\tilde{\partial}^{k-1}=M.$ (2.5)
[…]
3. We have two possible cases of such monomials:
@@
    $M=f_{i}(x)\tilde{A}_{i}{\tilde{\partial}^{\mathop{\rm\bf ord}(M)}}$. Notice that $\tilde{A}_{i}=e^{(\xi^{i}-1)x\tilde{\partial}}$ and $(\xi^{i}-1)\tilde{\partial}$ is the root of $L(t)$ with multiplicity 1. Then according to Lemma 2.8, such ODE has a special solution $H_{0}=xg_{i}(x)\tilde{A}_{i}{\tilde{\partial}^{\mathop{\rm\bf ord}(M)}}$ (with $deg_{x}(f_{i})=deg_{x}(g_{i})$).
@@
    $M=b_{j}\tilde{B}_{j}{\tilde{\partial}^{\mathop{\rm\bf ord}(M)}}$, where $b_{j}\in\tilde{K}$. Notice that $\tilde{B}_{j}=cx^{j-1}\tilde{\partial}^{j-1}\tilde{\delta}=cx^{j-1}\tilde{\partial}^{j-1}e^{-x\tilde{\partial}}$, and $\tilde{\partial}$ is not a root of $L(t)$. Then according to Lemma 2.8, we know the special solution of this ODE will be in the form
    $H_{0}=h_{j}(x)\tilde{\delta}{\tilde{\partial}^{\mathop{\rm\bf ord}(M)}}$
@@
[…]
Writing this solution in a canonical form and replacing $\tilde{\partial}$ by $D$, we get, as in the proof of propostion 2.1, the original solution $H$. Since $Sdeg_{A}(\sum_{i=0}^{k-1}c_{i}{A}_{k,i})=0$, and in view of lemma 2.6, we get our assertion. Thus we complete the proof.
@@
The condition about existence of $H$ from lemma is essential: for example $\partial^{4},H]=B_{10}\partial^{4}$ doesn’t have any HCPC solution. But we will see if $M$ doesn’t contain $B_{j}$ such $H$ must exist, see lemma 2.11.
@@
    $Sdeg_{A}(T)\leq Sdeg_{A}(H)+Sdeg_{A}(M)$ (here we assume $\infty+n=-\infty$).
@@
Consider the following 4 cases (below we assume in all formulae that $0^{0}:=1$):
@@
(here $\delta_{i}^{j-u}$ is the Kronecker delta), because if i=jui=j-u, then $i-1+u+v=j-1+v\geq 0$ and $B_{i}D^{u+v}\neq 0$.
2. $H=b_{i}B_{i}D^{u}$, where $i\geq 1-u$ if $u<0$, $M=a_{l,m}\Gamma_{m}A_{j}D^{v}$, $b_{i},a_{l,m}\in\tilde{K}$. By Lemma 2.4 item 9 we know $B_{i}D^{u}=D^{u}B_{i+u}$, hence $i\geq 1-u$, otherwise $H=0$. Then we have
@@
And by item 1 we know $B_{i}B_{j}=\delta^{i}_{j}B_{i}$, hence we know if $u-v+1\leq i\leq-u$, then
@@
But we already assume $i\geq 1-u$, hence $B_{i}D^{u}D^{v}=B_{i}D^{u+v}$, again by Lemma 2.4 item 9, we know it is 0 when i1+u+v<0i-1+u+v<0.
@@
where $\lambda=a_{l,m}b_{i}\xi^{j(i-v-1)}(i-v-1)^{m}$. This is because when iv<1i-v<1 we have $D^{v}B_{i}=B_{i-v}D^{v}=0$ (c.f. Lemma 2.4 item 9 and by assumptions for $r\leq 0$, $B_{r}=0$), so that
@@
The last equality holds obviously when $v\geq 0$ or $u\leq 0$. When $v<0,u>0$, notice that $i>0$, so that iv>vi-v>-v, hence for any $u-v+1\leq s\leq-v$, we have $B_{i-v}B_{s}=\delta_{s}^{i-v}=0$ by Lemma 2.4 item 5. Hence
@@ 1. Assume
Since $\partial^{k},A_{i}]=0$, we know by lemma 2.10 that
$\partial^{k},H_{i;m-k}A_{i}D^{m-k}]=M_{i;m}A_{i}D^{m}\qquad[\mod B_{j}],$ (2.10)
where $\mod B_{j}]$ denote denote terms containing $B_{j}$s. Now assume
@@
Since $\partial^{k},h_{0,i;m-k}A_{i}D^{m-k}]$ contains only terms with $B_{j}$, we can ignore terms with $h_{0,i;m-k}$ by looking for a solution. Assume
@@
Using Lemma 2.10, we can directly calculate $\partial^{k},H_{i;m-k}A_{i}D^{m-k}]$. Then we get the linear system for the unknown coefficients $h_{1,i;m-k},\dots,h_{t_{i},i;m-k}$}:
$\begin{pmatrix}k&*&\dots&*\\ 0&2k&\dots&*\\ \dots&\dots&\dots&\dots\\ 0&0&\dots&t_{i}k\end{pmatrix}\begin{pmatrix}h_{1,i;m-k}\\ h_{2,i;m-k}\\ \dots\\ h_{t_{i},i;m-k}\end{pmatrix}=\begin{pmatrix}m_{0,i;m}\\ m_{1,i;m}\\ \dots\\ m_{t_{i}-1,i;m}\end{pmatrix}$ (2.11)
@@
$\partial^{k},\tilde{H}]=M\quad[\mod B_{j}]$
Now assume
@@
with $H=\tilde{H}+\bar{H}$. The equation $\partial^{k},H]=M$ becomes
$\partial^{k},\bar{H}]=M-[\partial^{k},\tilde{H}$] (2.12)
@@
    $m\geq k$, in such case no $B_{j}$ can appear on the right hand side of (2.10). So we can simply put $h_{0,0;m-k},\dots,h_{0,k-1;m-k}$ all 0. We will get $\partial^{k},H]=M$.
@@
    where $b_{1-m},\dots,b_{k-m}$ are already known on the right hand side. Collecting this linear system, we find we have $k$ variables and $k$ equations, the coefficient matrix is a Vandermonde matrix (hence always of full rank). Thus we can always solve the equation and find the coefficients $h_{0,0;m-k},\dots,h_{0,k-1;m-k}$. After that we get $\partial^{k},\tilde{H}+\bar{H}]=M$.
@@
    $0<m<k$, in this case the same arguments as in case b) work, and we omit the details. We only need to notice that here we have $k$ variables but (km)(k-m) equations, and the coefficient matrix becomes the full rank sub-matrix of the Vandermonde matrix. So the solution exists but might be not unique.